# Improving generalisation via anchor multivariate analysis

### v2

@@ Abstract
We introduce a causal regularisation extension to anchor regression (AR) for improved out-of-distribution (OOD) generalisation. We present anchor-compatible losses, aligning with the anchor framework to ensure robustness against distribution shifts. Various multivariate analysis (MVA) algorithms, such as (Orthonormalized) PLS, RRR, and MLR, fall within the anchor framework. We observe that simple regularisation enhances robustness in OOD settings. Estimators for selected algorithms are provided, showcasing consistency and efficacy in synthetic and real-world climate science problems. The empirical validation highlights the versatility of anchor regularisation, emphasizing its compatibility with MVA approaches and its role in enhancing replicability while guarding against distribution shifts. The extended AR framework advances causal inference methodologies, addressing the need for reliable OOD generalisation.
Keywords Distribution shift  $\cdot$ Robust estimation  $\cdot$ Domain generalisation  $\cdot$ Causality  $\cdot$ Multivariate analysis  $\cdot$ Detection and attribution  $\cdot$ Climate Change
@@ 1 Introduction
Data sources in contemporary machine learning applications are often heterogeneous, leading to potential distribution shifts Sugiyama and Kawanabe (2012); Shen et al. (2021). This is a particularly relevant problem in computer vision Csurka (2017), healthcare Zhang et al. (2021), finance, Earth and climate sciences (Tuia et al., 2016; Kellenberger et al., 2021) and social sciences, as variations in data patterns can significantly impact model performance and generalisation in the out-of-distribution (OOD) setting, also referred to as domain generalisation (Shen et al., 2021; Zhou et al., 2023). Various frameworks have been proposed to formally address the emergence of distribution shifts during the testing phase Peters et al. (2016); Arjovsky et al. (2020). In cases where the data distribution is entailed by a Structural Causal Model (SCM) Peters et al. (2017), one can consider distribution shifts arising from intervention on specific variables of the SCM. Notably, the Instrumental Variable (IV) regression exhibits robustness to arbitrarily strong interventions (Bowden and Turkington, 1990).
However, pursuing algorithms robust to strong interventions may be overly conservative, especially when prior knowledge is available regarding the intervention strength that generates the distribution shift. Anchor Regression (AR) addresses this challenge by explicitly considering interventions on exogenous variables up to a specified strength (Rothenhäusler et al., 2018). This approach results in a straightforward and manageable regularisation of the classical Ordinary Least Squares (OLS) method, offering a solution for achieving distributionally robust linear regression.
@@
Several AR extensions have been proposed, demonstrating its versatility in handling different scenarios. Notable examples include (1) Kernel Anchor Regression (Shi and Xu, 2023), which extends AR to nonlinear SCMs by working on Reproducing Kernel Hilbert Spaces; (2) Targeted Anchor Regression Oberst et al. (2021), which exploits prior knowledge of the direction of intervention on the anchor variables; and (3) Proxy Anchor Regression Oberst et al. (2021), which addresses cases where the anchor is unobserved but proxy variables are available. Closer to our work, Kook et al. (2022) introduce a distributional extension of AR, expanding its application beyond the $\ell_{2}$ loss to accommodate censored or discrete targets.
All the literature above focuses on regression problems with a one-dimensional target variable and mainly employs least-squared error as the loss function. This paper extends the anchor framework introduced by Rothenhäusler et al. (2018) and contributes in two distinct ways. We first broaden its applicability to various loss functions, ensuring compatibility and robustness against distribution shifts caused by constrained interventions on the anchor variables. This extension redefines conventional MVA algoritms (Borchani et al., 2015; Arenas-Garcia et al., 2013), within the anchor framework, including algorithms such as Partial Least Squares (PLS), Orthonormalized PLS or Reduced Rank Regression (RRR) and Multilinear Regression (MLR). This signifies that their respective loss functions are anchor-compatible , and applying straightforward regularisation enhances their robustness in testing.
After reviewing the AR framework in §2, we study robustness through AR-based regularisation in §3 (anchor-compatible losses and their role in MVAs) and discuss the key links to causality and invariance (cf. §4). Then, we provide estimators for the selected algorithms and examine their consistency (cf. Section 5), supporting our theoretical findings through an extended set of simulation experiments. Finally, we showcase their efficacy in a challenging real-world climate science problem (cf. §6.2). We conclude in §7. The presented AR framework enhances causal inference methods by improving robust OOD generalisation. The code used in this work can be found at this GitHub repository.
@@ 2 Anchor framework
This section delineates the class of linear SCMs relevant to our work (§2.2) and offers insight into why anchor regularisation can be employed to obtain distributionally robust estimators (§2.3) for a more general class than the $\ell_{2}$ loss.
2.1 Notation
Bold capital letters $\mathbf{U}$ represent matrices, bold lowercase letters $\mathbf{u}$ denote vectors, regular lowercase letters $u$ signify scalars, and uppercase letters $U$ represent random variables (potentially multivariate) which for notation sakes we assume centred. Given two random vector $U$ and $V$ we define $UV$ as the row-wise stacking operation $UV=(U^{T},V^{T})^{T}$. The symbol $C_{UV}$ denotes Kronecker products involving random variables, i.e. $C_{UV}=UV\otimes UV$and to simplify the notation we write $C_{U}=U\otimes U$. Since we assume centred random variables $\Sigma_{UV}=\mathbb{E}[C_{UV}]$ and $\Sigma_{U}=\mathbb{E}[C_{U}]$. By a slight abuse of notation, we denote $C_{UV|A}=\mathbb{E}\left[UV|A\right]\otimes\mathbb{E}\left[UV|A\right]$ and $\Sigma_{UV|A}$ its expectation. [^1] Not to be confused with the conditional covariance matrix $Cov(UV|A)$. Our notation statisfies the total variance-covariance formula $\Sigma_{UV}=\mathbb{E}[Cov(UV|A)]+\Sigma_{UV|A}$. Again for simplification, we write $\sigma_{U}=\mathbb{E}[C_{U}]$ and $\sigma_{UV}=\mathbb{E}[U\otimes V]$. We denote the identity matrix with $\mathbf{I}$, and its dimension is implied by the context when not explicitly stated. Finally, .$\|.\|_{F}$ represents the Froebinius norm.
@@ 2.2 The Class of Anchor SCM
Figure 1: DAG studied in this work and induced by the SCM in (7).
In the following, we assume the Directed Acyclic Graph (DAG) in Fig. 1, where $X\in\mathbb{R}^{d}$ denotes the observed covariates, $Y\in\mathbb{R}^{p}$ represents the response (or target) variables, $H\in\mathbb{R}^{r}$ are unobserved variables potentially confounding the causal relation between $X$ and $Y$ and $A\in\mathbb{R}^{q}$ the exogenous (so-called anchor) variables. Following Rothenhäusler et al. (2018), we assume that the distribution of $(X,Y,H)$ is entailed by the following linear SCM $\mathcal{C}$:
@@
where $\mathbf{B}\in\mathbb{R}^{(d+p+r)\times(d+p+r)}$ and $\mathbf{M}\in\mathbb{R}^{(d+p+r)\times q}$ are unknown constant matrices and $\varepsilon\in\mathbb{R}^{d+p+r}$ is a vector of random noise. We assume that $A$ and $\varepsilon$ are independent and have finite variances, $\varepsilon$ components are independent, and $X$ and $Y$ have zero mean. Now, assuming that $(\mathbf{I}-\mathbf{B})$ is invertible, Eq. (7) can be written as:
$\begin{pmatrix}X\\ Y\\ H\end{pmatrix}=(\mathbf{I}-\mathbf{B})^{-1}(\varepsilon+\mathbf{M}A).$ (8)
The existence of $(\mathbf{I}-\mathbf{B})^{-1}$ is satisfied if the linear SCM is acyclic. Since our main interest lies in defining the variance-covariance $\Sigma_{XY}$ of $X$ and $Y$, we use $\mathbf{D}\in\mathbb{R}^{(d+p)\times(d+p+r)}$ to express:
@@
whose existence is implied by the existence of $(\mathbf{I}-\mathbf{B})$ inverse.
2.3 Anchor-Optimal Space
Let us now build upon (9) to illustrate how distributional robustness can be formulated in a causal context and give some intuition of the extension we propose in this work. Later, we will delve into the crucial role of $\Sigma_{XY}$ and its decomposition with $\varepsilon$ and anchor variables $A$. We focus on creating algorithms utilising $\Sigma_{XY}$. Note that, from (9), we can easily express $\Sigma_{XY}$ as
@@ 2.3 Anchor-Optimal Space
which combines the variance from observed exogenous (anchors) and unobserved exogenous (noise) variables.
Let us consider the triplet $(X,Y,H)\sim\operatorname{\mathbb{P}_{\nu}}$, where $\operatorname{\mathbb{P}_{\nu}}=\mathbb{P}^{do(A:=\nu)}_{\mathcal{C}}$ represents the perturbed distribution resulting from intervention (Peters et al., 2017) on the anchor variable $A$, where the distribution of $\nu$ is assumed to be mean-centred. From (10), the interventional variance-covariance matrix $\Sigma_{XY}^{do(A:=\nu)}$ can be expressed as
$\Sigma_{XY}^{do(A:=\nu)}$ $=\mathbf{D}\Sigma_{\varepsilon}\mathbf{D}^{T}+\mathbf{D}\mathbf{M}\Sigma_{\nu}\mathbf{M}^{T}\mathbf{D}^{T}.$ (11)
By imposing the constraint $\Sigma_{\nu}\preceq\gamma\Sigma_{A}$ [^2] Here, $U\preceq V$ means that $V-U$ is positive semi-definite. (the rationale for which will be developed in section §3.1), we can bound the interventional variance-covariance matrix $\Sigma_{XY}^{do(A:=\nu)}$ using the training covariances $\Sigma_{\varepsilon}$ and $\Sigma_{A}$. The anchor regularizer parameter $\gamma\in\mathbb{R}^{+}$, controls the amount of causal regularisation by constraining interventions on $A$. This yields the specific distribution class:
@@
As demonstrated in Eq. (31) and Eq. (30), both $\Sigma_{\varepsilon}$ and $\Sigma_{A}$ can be expressed in terms of $\Sigma_{XY|A}$ and $\Sigma_{XY}$ of the training distribution, which leads to the following inequality:
$\Sigma_{XY}^{do(A:=\nu)}\preceq\Sigma_{XY}+(\gamma-1)\Sigma_{XY|A}.$ (13)
Consequently, the interventional covariance matrix $\Sigma_{XY}^{do(A:=\nu)}$
can be bounded using only training data according to Eq. (13). As we will develop in Section 3.1, all algorithms exploiting $\Sigma_{XY}$ can thus be anchor-regularised, leading to distributionally robust estimators.
An example is the $\ell_{2}$ loss, defined for a one-dimensional response variable as $\mathbb{E}[\mathcal{L}(X,Y;\mathbf{b})]=\mathbb{E}[(Y-\mathbf{b}^{T}X)^{2}]=\Sigma_{Y}-2\mathbf{b}^{T}\sigma_{XY}+\mathbf{b}^{T}\Sigma_{X}\mathbf{b}$, which is evidently linear on $\Sigma_{XY}$. Similarly, the PLS algorithm Abdi (2010) aims to find a pair of projection matrices $\mathbf{W}_{x}\in\mathbb{R}^{d\times u}$ and $\mathbf{W}_{y}\in\mathbb{R}^{p\times v}$ that maximise the covariance between the transformed data $X\mathbf{W}_{x}$ and $Y\mathbf{W}_{y}$, i.e. assuming centered $X$ and $Y$, we aim to maximise $\text{tr}(\mathbf{W}_{x}^{T}\Sigma_{XY}\mathbf{W}_{y})$ while constraining $\mathbf{W}_{x}$ and $\mathbf{W}_{y}$ to be orthogonal. Here $u$ and $v$ represent the dimension of the projected spaces. In this case, as the loss is also linear on matrix $\Sigma_{XY}$, i.e. PLS can be anchor-regularised, cf. §3.1.
An intuitive way of thinking about anchor regularisation is to consider that by perturbing the training data (see Eq. (21)), one obtains a dataset that is close to the worst-case scenario we could observe during testing. These observations leads us to consistently define a class of loss functions and demonstrate their distributional robustness properties over the class of perturbed distributions $C^{\gamma}$ defined in (12).

### v3

# Out-of-distribution robustness for multivariate analysis via causal regularisation

@@ Abstract
We propose a regularisation strategy of classical machine learning algorithms rooted in causality that ensures robustness against distribution shifts. Building upon the anchor regression framework, we demonstrate how incorporating a straightforward regularisation term into the loss function of classical multivariate analysis algorithms, such as (orthonormalized) partial least squares, reduced-rank regression, and multiple linear regression, enables out-of-distribution generalisation. Our framework allows users to efficiently verify the compatibility of a loss function with the regularisation strategy. Estimators for selected algorithms are provided, showcasing consistency and efficacy in synthetic and real-world climate science problems. The empirical validation highlights the versatility of anchor regularisation, emphasizing its compatibility with multivariate analysis approaches and its role in enhancing replicability while guarding against distribution shifts. The extended anchor framework advances causal inference methodologies, addressing the need for reliable out-of-distribution generalisation.
@@ 1 INTRODUCTION
Data sources in contemporary machine learning applications are often heterogeneous, leading to potential distribution shifts (Sugiyama and Kawanabe, 2012; Liu et al., 2023). This is a particularly relevant problem in computer vision (Csurka, 2017), healthcare (Zhang et al., 2021), Earth and climate sciences (Tuia et al., 2016; Kellenberger et al., 2021), and social sciences (Jin et al., 2023), as variations in data patterns can significantly impact model performance and generalisation in the out-of-distribution (OOD) setting, also referred to as domain generalisation (Liu et al., 2023; Zhou et al., 2023).
In this work, we address the problem of predicting a target $Y\in\mathcal{Y}$ from covariates $X\in\mathcal{X}$ where the training data consists of samples from a subset $\mathcal{Q}$ of a broader class of distributions $\mathbb{Q}$. Our objective is to achieve strong performance across the entire class $\mathbb{Q}$. This challenge can be framed as the following minimax problem:
where $\mathcal{L}(X,Y;\mathbf{\Theta})$ is a loss function over $X$ and $Y$ with parameters $\mathbf{\Theta}$. The choice of the class $\mathbb{Q}$ is particularly important as different classes of distributions lead to various types of robustness. The class $\mathbb{Q}$, often referred to as the uncertainty set, can be constrained using distributional metrics such as $f$-divergence (see, e.g., Namkoong and Duchi, 2016) or the Wasserstein distance (see, e.g., Esfahani and Kuhn, 2015). The challenge of OOD generalisation is also closely related to the field of transfer learning (see, e.g., Zhuang et al., 2021). For an in-depth review of OOD generalisation, we recommend consulting (Liu et al., 2023; Zhou et al., 2023).
Of particular interest for this work, causal inference can be understood as a distributionally robust estimation where the class $\mathbb{Q}$ is the class of all distributions arising from arbitrarily large interventions on the variables of an SCM (see Meinshausen, 2018; Liu et al., 2023). Notably, the Instrumental Variable (IV) regression exhibits robustness to arbitrarily strong interventions on the instruments under some structural assumptions (Bowden and Turkington, 1990). However, it might be overconservative to pursue algorithms robust to arbitrarily strong interventions, especially when assuming access to knowledge regarding the intervention strength. Rothenhäusler et al. (2018) address this challenge by constraining the intervention to exogenous (so-called anchor) variables and by bounding the intervention strength. This approach, known as Anchor Regression (AR), results in a straightforward regularisation of the Ordinary Least Squares (OLS) algorithm, enabling robust linear regression.
More concretely, in its original form, AR can be written, in the population case, as
$b^{\gamma}=\operatorname{\text{argmin}}_{\gamma}$ $\mathbb{E}\left[((I-P_{A})(Y-X^{T}b))^{2}\right]$
$+\gamma\mathbb{E}\left[(P_{A}(Y-X^{T}b))^{2}\right],$
with $Y\in\mathbb{R}$. The second term acts as a regulariser that, for $\gamma>1$, enforces the minimization of the residuals’ projection onto the linear space defined by $A$. This term is akin to the two-stage least squares used in IV regression, as it promotes decorrelation between the anchor and the residuals, thereby making the model more robust to distribution shifts arising from interventions on the anchor. This regularisation approach has proven useful in various applications, notably in environmental fields where Oberst et al. (2021) applied it to air-quality prediction, and Sippel et al. (2021) and Székely et al. (2022) utilized it for the Detection and Attribution (D&A) of climate change.
@@
Several AR extensions have been proposed, demonstrating the theoretical versatility of this framework. Notable examples include Kernel Anchor Regression (Shi and Xu, 2023), which extends AR to nonlinear SCMs by working on reproducing kernel Hilbert spaces; targeted anchor regression (Oberst et al., 2021), which exploits prior knowledge of the direction of intervention on the anchor variables; and proxy anchor regression (Oberst et al., 2021), which addresses cases where the anchor is unobserved but proxy variables are available. In Kook et al. (2022), authors introduce a distributional extension of AR, expanding beyond the $\ell_{2}$ loss to accommodate censored or discrete targets.
Previous literature has focused on regression problems with one-dimensional target variables and ordinary least squares as the loss function. Building upon the original approach of AR, we introduce a regularisation strategy applicable to a broader class of algorithms, including some classical MultiVariate Analysis (MVA) techniques (Bilodeau and Brenner, 1999; Arenas-Garcia et al., 2013; Borchani et al., 2015), also known as multi-output algorithms. This contributes on three fronts. First, we extend the applicability of AR algorithms to various loss functions, thereby ensuring compatibility and robustness against distribution shifts induced by constrained interventions on anchor variables. Second, we redefine conventional MVA algorithms within the anchor framework, encompassing techniques such as Partial Least Squares (PLS), Orthonormalized PLS (OPLS), Reduced Rank Regression (RRR), and Multilinear Regression (MLR). This adaptation ensures that their respective loss functions are anchor-compatible, and the application of straightforward regularisation provides robustness during testing. This, allows us to leverage the power and versatility of a diverse array of MVA algorithms while providing robustness guarantees against distributional shifts. Third, by introducing a new class of anchor-compatible loss functions, we offer a simple and straightforward method to determine whether an algorithm could benefit from anchor regularization and its generalization properties.
The code used in this work can be found on GitHub.
@@ 2 ANCHOR FRAMEWORK
For notation, bold capital letters $\mathbf{U}$ represent matrices, bold lowercase letters $\mathbf{u}$ denote vectors, regular lowercase letters $u$ signify scalars, and uppercase letters $U$ represent random variables (potentially multivariate) which we assume are centered. Given two random vector $U$ and $V$ we define $UV$ as the row-wise stacking operation $UV=(U^{T},V^{T})^{T}$. The symbol $C_{UV}$ denotes Kronecker products involving random variables, i.e. $C_{UV}=UV\otimes UV$ and to simplify the notation we write $C_{U}=U\otimes U$. Since we assume centred random variables $\Sigma_{UV}=\mathbb{E}[C_{UV}]$ and $\Sigma_{U}=\mathbb{E}[C_{U}]$. By a slight abuse of notation, we denote $C_{UV|A}=\mathbb{E}\left[UV|A\right]\otimes\mathbb{E}\left[UV|A\right]$ and $\Sigma_{UV|A}$ its expectation. [^1] Not to be confused with the conditional covariance matrix $\mathrm{Cov}(UV|A)$. Our notation satisfies the total variance-covariance formula $\Sigma_{UV}=\mathbb{E}[\mathrm{Cov}(UV|A)]+\Sigma_{UV|A}$. Again for simplification, we write $\sigma_{U}=\mathbb{E}[C_{U}]$ and $\sigma_{UV}=\mathbb{E}[U\otimes V]$. We denote the identity matrix with $\mathbf{I}$, and its dimension is implied by the context when not explicitly stated. Finally, .$\|.\|_{F}$ represents the Froebinius norm.
@@ 2.1 The Class of Anchor SCM
Figure 1: Directed Acyclic Graph (DAG) analyzed in this work, as induced by the Structural Causal Model (SCM) described in (2). The directions of the arrows between $X$, $Y$, and $H$ are flexible, provided that the graph remains acyclic. All possible configurations of the DAG can be seen in Fig. 4 in Appendix.
In the following, we assume the Directed Acyclic Graph (DAG) in Fig. 1, where $X\in\mathbb{R}^{d}$ denotes the observed covariates, $Y\in\mathbb{R}^{p}$ represents the response (or target) variables, $H\in\mathbb{R}^{r}$ are unobserved variables potentially confounding the causal relation between $X$ and $Y$ and $A\in\mathbb{R}^{q}$ the exogenous (so-called anchor) variables. We assume that the distribution of $(X,Y,H)$ is entailed by the following linear SCM $\mathcal{C}$:
@@
where $\mathbf{B}\in\mathbb{R}^{(d+p+r)\times(d+p+r)}$ and $\mathbf{M}\in\mathbb{R}^{(d+p+r)\times q}$ are unknown constant matrices and $\varepsilon\in\mathbb{R}^{d+p+r}$ is a vector of random noise. We assume that $A$ and $\varepsilon$ are independent and have finite variances, $\varepsilon$ components are independent, and $X$ and $Y$ have zero mean. The model in Eq. 2 generalises the Instrumental Variables setting, offering a flexible framework that encompasses any linear model—whether or not hidden confounders are present—where an exogenous variable is observed, without requiring an exclusion restriction. Assuming that $(\mathbf{I}-\mathbf{B})$ is invertible, which is satisfied if the linear SCM is acyclic, we can easily express
@@
where $\mathbf{D}\in\mathbb{R}^{(d+p)\times(d+p+r)}$ existence is implied by the existence of $(\mathbf{I}-\mathbf{B})^{-1}$.
2.2 Bound on the Perturbed Covariance
Let us now build upon Eq. (3) to illustrate how distributional robustness can be formulated in a causal context and give some intuition of the proposed regularisation strategy. We show how anchor regularisation can be formulated in terms of covariances $\Sigma_{XY}$ and $\Sigma_{XY|A}$, instead of squared residuals as originally proposed in Rothenhäusler et al. (2018), the former being more general. This relies on the simple idea that the squared residuals loss is a linear form on $\Sigma_{XY}$ and anchor regularisation simply adds a constraint on the perturbed covariance matrix $\Sigma_{A}$ (see Eq. (5)).
Note that, from Eq. (3), we can easily express $\Sigma_{XY}$ as
@@ 2.2 Bound on the Perturbed Covariance
which combines the variance from exogenous observed (anchors) and unobserved (noise) variables.
Consider the triplet $(X,Y,H)\sim\operatorname{\mathbb{P}_{\nu}}$, where $\operatorname{\mathbb{P}_{\nu}}=\mathbb{P}^{do(A\sim\nu)}_{\mathcal{C}}$ represents the perturbed distribution resulting from intervention (see Peters et al., 2017, Sec. 3.2) on the anchor variable $A$, where the distribution of $\nu$ is assumed to be mean-centered. From Eq. (4), the perturbed variance-covariance matrix $\Sigma_{XY}^{do(A\sim\nu)}$ can be expressed as
$\Sigma_{XY}^{do(A\sim\nu)}=\mathbf{D}\Sigma_{\varepsilon}\mathbf{D}^{T}+\mathbf{D}\mathbf{M}\Sigma_{\nu}\mathbf{M}^{T}\mathbf{D}^{T},$
with $\Sigma_{\nu}=\mathbb{E}[\nu\nu^{\top}]$. By imposing the constraint $\Sigma_{\nu}\preceq\gamma\Sigma_{A}$ [^2] Here, $U\preceq V$ means that $V-U$ is positive semi-definite. , we can bound the interventional variance-covariance matrix $\Sigma_{XY}^{do(A\sim\nu)}$ using the training covariances $\Sigma_{\varepsilon}$ and $\Sigma_{A}$. An intuition for this constraint is that during testing, the covariance matrix of the anchor variable can be expected to be scaled by a factor of at most $\gamma$ in all directions, thus constraining the strength of the intervention. The anchor regulariser parameter $\gamma\in\mathbb{R}^{+}$ thus controls the amount of causal regularisation. This yields the specific distribution class:
@@
As demonstrated in Eq. (20) and Eq. (19) in Appendix, both $\Sigma_{\varepsilon}$ and $\Sigma_{A}$ can be expressed in terms of $\Sigma_{XY|A}$ and $\Sigma_{XY}$ of the training distribution, which leads to the following inequality:
$\Sigma_{XY}^{do(A\sim\nu)}\preceq\Sigma_{XY}+(\gamma-1)\Sigma_{XY|A}.$ (6)
Consequently, the interventional covariance matrix $\Sigma_{XY}^{do(A\sim\nu)}$ can be bounded using only the training (or unperturbed) distribution according to Eq. (6). As we will develop in §3.1, algorithms exploiting $\Sigma_{XY}$ can thus be anchor-regularised, leading to distributionally robust estimators. An example is the $\ell_{2}$ loss, defined for a one-dimensional response variable as $\mathbb{E}[\mathcal{L}(X,Y;\mathbf{b})]=\mathbb{E}[(Y-\mathbf{b}^{T}X)^{2}]=\sigma_{Y}-2\mathbf{b}^{T}\sigma_{XY}+\mathbf{b}^{T}\Sigma_{X}\mathbf{b}$, which is evidently linear on $\Sigma_{XY}$. Similarly, the PLS algorithm (Abdi, 2010) aims to find a pair of matrices $\mathbf{W}_{x}\in\mathbb{R}^{d\times u}$ and $\mathbf{W}_{y}\in\mathbb{R}^{p\times v}$ that maximise the covariance between the transformed data $X\mathbf{W}_{x}$ and $Y\mathbf{W}_{y}$, i.e. assuming centered $X$ and $Y$, we aim to maximise $\text{tr}(\mathbf{W}_{x}^{T}\Sigma_{XY}\mathbf{W}_{y})$ while constraining the columns of $\mathbf{W}_{x}$ and $\mathbf{W}_{y}$ to each be orthogonal. In this case, as the loss is also linear on matrix $\Sigma_{XY}$, i.e. PLS can be anchor-regularised, cf. §3.1. These observations lead us to define a class of loss functions and demonstrate their robustness over the perturbed distributions $C^{\gamma}$ defined in Eq. (5).